### v2

# HumanSense: From Multimodal Perception to Empathetic Context-Aware Responses through Reasoning MLLMs

Abstract
While Multimodal Large Language Models (MLLMs) show immense promise for achieving truly human-like interactions, progress is hindered by the lack of fine-grained evaluation frameworks for human-centered scenarios, encompassing both the understanding of complex human intentions and the provision of empathetic, context-aware responses. Here we introduce HumanSense, a comprehensive benchmark designed to evaluate the human-centered perception and interaction capabilities of MLLMs, with a particular focus on deep understanding of extended multimodal contexts and the formulation of rational feedback. Our evaluation reveals that leading MLLMs still have considerable room for improvement, particularly for advanced interaction-oriented tasks. Supplementing visual input with audio and text information yields substantial improvements, and Omni-modal models show advantages on these tasks. Furthermore, we argue that appropriate feedback stems from a contextual analysis of the interlocutor’s needs and emotions, with reasoning ability serving as the key to unlocking it. Accordingly, we employ a multi-stage, modality-progressive reinforcement learning to enhance the reasoning abilities of an Omni model, achieving substantial gains on evaluation results. Additionally, we observe that successful reasoning processes exhibit highly consistent thought patterns. By designing corresponding prompts, we also enhance the performance of non-reasoning models in a training-free manner. Project page: https://github.com/antgroup/HumanSense
footnotetext: Co-first author. Project lead. Corresponding author. 1National Key Laboratory of Human-Machine Hybrid Augmented Intelligence, National Engineering Research Center for Visual Information and Applications, Institute of Artificial Intelligence and Robotics, Xi’an Jiaotong University. 2Ant Group.
Introduction
Science fiction (Quantic Dream 2018; her 2013) often portrays a future where artificial intelligence serves not merely as a tool for task execution, but as a human companion offering social support and emotional connection. The fundamental evolution from narrow, task-oriented systems to Artificial General Intelligence is predicated on the capacity to comprehend human intentions from speech, expressions, and body language, thereby enabling appropriate responses.
Multimodal Large Language Models (MLLMs) (Xu et al. 2025; Hurst et al. 2024; Anthropic 2024; Team et al. 2023) represent a promising pathway toward realizing this vision. Their ability to holistically process visual, auditory, and textual information enables a comprehensive understanding of users and environments. MLLMs also have the potential to deeply analyze perceived information (Guo et al. 2025) and subsequently plan appropriate feedback, which is not limited to textual responses, but can include suitable emotions, tones, and gesture labels in temporal sequences. Such outputs can be further integrated with video generation (Meng et al. 2025; Qin et al. 2025), speech synthesis (Wang et al. 2023; Suno AI 2023), and talking head (Li et al. 2024c, b; Zhang et al. 2024c; Zheng et al. 2021) methods to provide a highly anthropomorphic interactive experience.
Achieving this goal first requires defining the necessary capabilities, evaluating model performance, and then applying optimization. However, existing benchmarks (Qi et al. 2025; Lin et al. 2024; Hu et al. 2025) lack targeted, fine-grained evaluation for these human-centered scenarios. To address this gap, we first define the primary capabilities needed for MLLMs in such scenarios: 1) multi-modal perception, 2) contextual understanding of implicit information, and 3) appropriate responses in multi-turn interactions. For interactive scenarios, we consider both response content and response strategies.
Figure 1: Demonstration of tasks in HumanSense. Note: The questions above are slightly simplified for clarity and brevity.
Based on the above considerations, we propose the HumanSense benchmark. This benchmark comprises 15 progressively challenging tests, totaling 3,882 questions derived from real-world records. In interactive tests, MLLMs are tasked with assuming the role of one party in the interaction and generating responses, which are then compared to real human records. We conduct a comprehensive evaluation of current leading MLLMs, including Vision-Language Models (Zhu et al. 2025; Team 2025), Omni models (Hurst et al. 2024; Xu et al. 2025), and Audio-Language Models (Chu et al. 2023). The results reveal significant room for improvement in human-centered scenarios, particularly in advanced interaction-oriented tasks. Modality ablation studies show that visual, auditory, and textual information all play crucial roles in high-level tasks, and omni models capable of jointly processing audio, video, and text exhibit a clear advantage.
Building on these findings, we propose that omni-modal reasoning can enhance the cognitive and interactive capabilities of MLLMs. This assertion stems from our observation that appropriate feedback in communication relies on thorough consideration of omni-modal context, the interlocutor’s needs, emotions, and personal characteristics. Accordingly, we employ a multi-stage, omni-modal reinforcement learning approach to build a reasoning omni model, resulting in substantial improvements in evaluations. Furthermore, we observe that successful reasoning processes exhibit highly consistent patterns. By designing corresponding prompts, we also enhance the performance of non-reasoning models in a training-free manner.
The path to Artificial General Intelligence requires long-term, multifaceted exploration. With this work, we aim to inspire the community to explore the potential of Omni MLLMs for improving human-centered AI interactions and helps shape this emerging direction.
Related Works
Multimodal Large Language Models.
LLMs (Bai et al. 2023; Mann et al. 2020; Radford et al. 2018, 2019; Touvron et al. 2023) have been extensively adopted for human behavior and sentiment analysis, facilitating applications including dialogue simulation, behavior prediction, and textual sentiment classification. These capabilities support diverse scenarios ranging from social media monitoring to automated customer support systems. However, textual information alone is often insufficient, as these models lack support for visual cues such as facial expressions and body language, which are essential for comprehensive human behavior analysis. Visual MLLMs (Lin et al. 2023; Chen et al. 2024; Li et al. 2024a; Wang et al. 2024b; Team et al. 2024; Yao et al. 2024) demonstrate strong capabilities in visual understanding, accurately recognizing emotions and behaviors through analysis of facial expressions and body language. However, a critical limitation of these models is their inability to process audio information, resulting in the loss of crucial auditory cues such as dialogue content, vocal intonations, and ambient sounds. This limitation creates significant gaps and introduces biases in their understanding of complex real-world scenarios. Omni models (Xu et al. 2025; Liu et al. 2025; Zhang et al. 2024b; Fu et al. 2025b; Fang et al. 2024; Zhang et al. 2023; Hurst et al. 2024), by contrast, integrate multiple modalities—including vision, language, and audio—to provide comprehensive simulation of complex human interactions. These models can process dialogue content while simultaneously analyzing visual cues, enabling more nuanced and accurate understanding of human communication dynamics.
MLLM Benchmarks.
With the advancement of multimodal large models, several evaluation benchmarks (Chen et al. 2024; Fu et al. 2025a; Wang et al. 2024c; Li et al. 2024d; Zhang et al. 2024a) have emerged, most of which focus on assessing video understanding capabilities. Additionally, StreamingBench (Lin et al. 2024) specializes in streaming video understanding. However, few benchmarks evaluate large models from a human-centered perspective, which is crucial for the practical deployment of such models in real-world scenarios. HumanOmniV2 (Yang et al. 2025) focuses on deciphering intentions, interpreting emotions, and detecting potential deception in an omni-modal manner. While HumanOmniV2 offers valuable insights into human-centered video understanding, it lacks the evaluation of the response planning or interactive capabilities.
Figure 2: Left: The structure and counts of evaluation tasks in HumanSense. Right: Benchmark statistics on video length.
HumanSense-Bench
Overview
We aim to systematically assess human-centered capabilities of MLLM through HumanSense framework: 1) Human-centered multi-modal perception, 2) Contextual understanding of implicit information, and 3) Response strategy in interactive scenarios, as Figure 1. The evaluation tasks are organized into a four-tier pyramid structure (L1–L4) according to increasing levels of difficulty, as shown in Figure 2 (Left):
  • Perception (L1 & L2): The base layer focuses on unimodal, surface perception tasks in L1, and L2 addresses multi-modal, long-duration complex perception tasks. Together, they form the foundational capabilities of intelligence.
  • Understanding (L3): Built upon perception, this layer evaluates whether the model can uncover implicit information embedded in conversations.
  • Response (L4): As the pinnacle of capabilities, this layer assesses the model’s ability to generate appropriate and rational responses in various interactive scenarios.
This design ensures the systematic nature of the evaluation: the model must first possess a solid perceptual foundation before advancing to deep understanding, ultimately enabling it to make informed feedback decisions at the top level. See Figure 2 for an overview of HumanSense tasks and dataset statistics on task quantity distribution and video length.
Task Definition
In human communication, diverse information is conveyed through different modalities. For example, visual expressions such as facial expressions and gestures can transmit emotional or semantic information; sound can directly express content or indirectly convey emotions. Perceiving these fundamental pieces of information is essential for interaction. We design the following multi-modal perception tasks.
L1: Surface Perception
  • Person Appearance (PA) evaluates the model’s fine-grained perception of facial appearance, as appearance constitutes a fundamental aspect of person identification. We leverage the data annotations from CelebV-HQ (Zhu et al. 2022) to design a series of multiple-choice questions. Each question asks whether the person in the video possesses a specific attribute, such as “Male”, “Young”, “Chubby”, “Rosy cheeks”, “Oval face”, or “straight hair”.
  • Action Recognition (AR) aims to evaluate the model’s ability to recognize atomic human actions. We ask MLLMs to identify the current action from a movie clip which comes from the AVA (Atomic Visual Actions) (Gu et al. 2018) dataset. The actions include individual behaviors (e.g., “walk”, “sleep”), interactions with objects (e.g., “Open a window”, “Row a boat”), and interactions between people (e.g., “Kiss a person”, “talk to a person”).
  • Gesture Recognition (GR) aims to evaluate the model’s ability to recognize hand gestures, which convey rich semantic information during communication. We construct single-choice questions using Jester (Materzynska et al. 2019) dataset, with gestures such as “Rolling Hand Forward” and “Sliding Two Fingers Up”.
  • Emotion Recognition (ER) examines the recognition of facial expressions, as they are the primary means of conveying emotions. Based on CelebV-HQ (Zhu et al. 2022) dataset, we generate single-choice questions using its built-in labels, asking the model to identify the emotion in videos such as “Happy”,“Sad”, “Disgust”, “Anger”, etc.
L2: Complex Perception
  • Complex Motion (CM) examines the description of extended complex action sequences, which relates to the understanding of target behaviors. We utilize the captions from ActivityNet dataset (Caba Heilbron et al. 2015) to construct single-choice questions, where the model must identify the correct description of the action performed in the long video clip.
  • Abnormal Classification (AC) evaluates the detection of abnormal human behaviors. We formulate single-choice questions based on the UCF-Crime100 dataset (Sultani, Chen, and Shah 2018). The abnormal events include “Stealing”, “Robbery”, or “Fighting”, etc.
  • Human Counting with ReID (HC) evaluates the model’s ability to recognize and remember individuals. We use the tracking dataset TAO (Dave et al. 2020) to calculate the number of individuals and ask the model about the total count of distinct persons appearing throughout a video. Some challenging questions involve camera transitions and the intermittent appearance and disappearance of humans.
  • Talking Person Recognition (TR) evaluates the ability to make judgments by integrating visually and auditorily perceived information. Based on RealTalk dataset (Geng et al. 2023), we extract video clips and formulate single-choice questions, where the model is required to identify which person is speaking a specific content.
The L1 and L2 tasks assess the perceptual abilities of MLLMs in relation to “seeing” and “hearing”. Achieving harmonious communication requires deep thinking about contextual content and providing appropriate responses that correspond to the interlocutor’s emotions. Accordingly, L3 examines the model’s capacity to “understand” implicit information during interactions, while L4 evaluates the model’s “response” abilities across different scenarios.
L3: Contextual Understanding
  • Familiarity Recognition (FR1) evaluates the model’s ability to understand human interpersonal interactions and perceive interpersonal closeness. Based on the duration of acquaintance between conversational participants in The Skin Deep channel, we construct single-choice questions to have the model determine the familiarity level between individuals in the videos.
  • Rapport Recognition (RR1) evaluates the model’s ability to perceive whether the communication atmosphere is harmonious. We annotate video chat content from The Skin Deep channel across multiple dimensions, including interaction frequency, communication atmosphere, and degree of viewpoint conflict, to construct single-choice questions. The evaluation encompasses whether the interactive atmosphere is pleasant and whether conflicts exist in the conversational content.
  • Relation Recognition (RR2) evaluates the model’s ability to determine human relationships through multi-modal context. Based on the relationships between conversational participants in The Skin Deep channel, we construct single-choice questions to have the model predict the relationship type between individuals in the videos, such as “Married”,“Siblings”,“Son/Mother”, etc.. The model must integrate information from visual appearance, age differences, and dialogue content to make a judgment.
  • Lie Detection (LD) evaluates the model’s ability to detect human lying. We formulated single-choice questions based on the SEUMLD dataset. The model is required to determine whether the speaker is lying in each video segment through analysis of visual and audio cues.
  • Fraud Recognition (FR2) evaluates the detection of specific fraudulent behaviors through approximately one-minute phone recordings, derived from the Telecom Fraud Texts dataset (Li, Zhang, and Jiang 2024). The fraud types include “Loan and credit card agency fraud”, “Impersonation of public security, judiciary, and government agencies” and “Impersonation of leaders or acquaintances”, etc.
L4: Feedback Strategy
  • Emo Strategy (ES) evaluates the MLLM’s ability to provide appropriate facial expression feedback during communication, creating an empathetic interaction experience. Based on the RealTalk dataset (Geng et al. 2023), we extract video clips using the speaker as input for questions, and annotated the listener’s facial expressions as answers to construct single-choice questions.
  • Psychological Chat (PC) evaluates the ability of MLLMs to generate appropriate responses in complex, long-context interactions. Here we use a professional online psychological dataset, Emotional First Aid dataset (Hai Liang Wang 2020), from which we constructed single-choice questions. The model is expected to select an appropriate response based on the previous multi-turn dialogue.
Data Construction
Question-Answer Generation.
HumanSense consists of 3,291 video-based questions and 591 audio-based questions. The related data is sourced from existing open-source datasets and YouTube videos. We construct Question-Answer (QA) pairs using templates, leveraging annotations and built-in labels from the existing datasets. We also use various off-the-shelf modules, including emotion recognition, Large Language Models(LLMs), and Optical Character Recognition (OCR), to analyze source data and extract task-relevant information. Detailed information on the construction for each task is available in the supplementary materials.
Models Avg. AR ER GR PA AC CM HC TR FR1 FR2 LD RR1 RR2 ES PC
L1 L2 L3 L4
HumanSense (tiny) Perf.
Human Level 0.875 0.917 0.933 0.767 0.933 0.967 0.967 0.889 0.900 0.900 0.800 0.533 0.967 0.833 0.867 0.933
GPT-4o 0.552 0.583 0.233 0.700 0.517 0.733 0.767 0.522 0.400 0.833 - 0.300 0.467 0.467 0.667 -
InternVL3-8B 0.558 0.417 0.467 0.533 0.433 0.833 0.767 0.567 0.333 0.733 - 0.667 0.433 0.433 0.633 -
Qwen2.5-Omni-7B 0.578 0.467 0.500 0.300 0.383 0.633 0.800 0.467 0.600 0.733 0.700 0.567 0.767 0.600 0.700 0.367
Qwen2-Audio-7B - - - - - - - - - - 0.333 - - - - 0.333
HumanSense Perf.
Proprietary Models (API)
GPT-4o 0.557 0.548 0.282 0.620 0.517 0.750 0.776 0.536 0.570 0.735 - 0.310 0.480 0.535 0.587 -
Vl-Model
LLaVA-Next-Video-7B 0.479 0.500 0.480 0.263 0.583 0.440 0.413 0.264 0.487 0.665 - 0.505 0.560 0.500 0.577 -
Qwen2-VL-7B 0.507 0.473 0.470 0.307 0.322 0.600 0.591 0.424 0.537 0.665 - 0.515 0.627 0.495 0.570 -
Qwen2.5-VL-7B 0.512 0.540 0.497 0.267 0.448 0.550 0.644 0.461 0.523 0.480 - 0.545 0.627 0.485 0.590 -
VideoLLaMA3-7B 0.520 0.543 0.463 0.323 0.517 0.530 0.694 0.561 0.493 0.610 - 0.515 0.587 0.400 0.530 -
LLaVA-OneVision-7B 0.521 0.545 0.510 0.400 0.592 0.620 0.676 0.268 0.503 0.600 - 0.530 0.560 0.430 0.543 -
InternVL3-8B 0.561 0.393 0.483 0.387 0.547 0.670 0.751 0.630 0.567 0.735 - 0.555 0.527 0.490 0.557 -
Audio-Model
Qwen2-Audio-7B - - - - - - - - - - 0.437 - - - - 0.399
Omni-Model
Ola-7B 0.525 0.557 0.463 0.263 0.573 0.320 0.420 0.371 0.597 0.785 0.733 0.565 0.653 0.640 0.577 0.567
IXC2.5-OmniLive-7B 0.544 0.508 0.467 0.257 0.338 0.660 0.584 0.544 0.533 0.780 0.415 0.500 0.560 0.470 0.463 0.324
Qwen2.5-Omni-7B 0.554 0.473 0.513 0.303 0.350 0.600 0.630 0.438 0.600 0.770 0.740 0.550 0.713 0.650 0.607 0.399
Table 1: Evaluation on HumanSense. Indicates results on the HumanSense (tiny) set, for comparison with human-level performance. While GPT-4o is designed as an omni-modal, the absence of audio input support in its current API precluded its evaluation on audio-related tasks (FR2, PC). For a fair comparison across all models, the overall average scores (Avg.) for omni-modals are calculated without the results from the two audio tasks. The best results for each metric are bolded.
Question-Answer Augmentation.
To improve evaluation generalizability and avoid evaluation bias, we augment both questions and answers. For each questions, we design diverse candidate templates for random selection. For answers, we balance the distribution of correct options and equalize the lengths of correct and incorrect choices. This QA augmentation also avoid reward hacking (Weng 2024) during the following Reinforcement Learning experiments.
Quality Control.
We conduct manual inspection on a twenty percent sample of all QA pairs, with particular focus on L3 and L4 tasks. For example, in the “Psychological Chat”, we ensure that context contains sufficient multi-turn dialogue. We also rigorously validate the correct options to ensure they provide professional and appropriate advice. In the“Rapport Recognition” task, we design scoring dimensions including interaction frequency, communication atmosphere, and content harmony for LLMs evaluation, followed by manual quality inspection of the results.
Figure 3: Performance Radar Charts on HumanSense (mini). The results include human-level performance with several state-of-the-art multimodal models.
Evaluation on HumanSense
We conduct a comprehensive evaluation of leading Multimodal Large Language Models (MLLMs) with sizes up to 10B, including: (1) Visual LLMs, which represent the most mainstream branch of MLLMs today; (2) Audio LLMs; and (3) Omni-modal LLMs that are natively designed for integrating vision, audio, and text. For Visual LLMs, we assess models such as Qwen2.5-VL (Team 2025), Qwen2-VL (Wang et al. 2024a), LLaVA-OneVision (Li et al. 2024a), LLaVANeXT-Video (Liu et al. 2024), VideoLLaMA3 (Zhang et al. 2025) and InternVL3 (Zhu et al. 2025). In the omni-modal models, we test Qwen2.5-Omni (Xu et al. 2025), IXC2.5OmniLive (Zhang et al. 2024b) and Ola (Liu et al. 2025). For audio LLMs, we evaluate Qwen2-Audio (Chu et al. 2023). Additionally, we test the powerful omni-model GPT-4o (Hurst et al. 2024). All evaluations are conducted in a zero-shot setting, employing the default prompts provided for each model. For video processing, we adhere to the respective official configurations for key parameters, including frame extraction methods, frames per second (FPS), and resolution. We will release our data and code to the community, aiming to facilitate broader evaluation across a diverse range of models.
To accommodate the characteristics of each model, Visual LLMs are exclusively evaluated on video tasks. Similarly, Audio LLMs are assessed solely on audio tasks. In contrast, omni-models, which are designed for multimodal processing, are required to complete all tasks.
Human Level Performance.
To establish a human performance benchmark, we curate a new evaluation set, dubbed HumanSense (tiny), comprising 450 questions randomly sampled across our tasks (30 per task). We first establish a human baseline by having evaluators independently answer each question. Subsequently, we compare the performance of several leading models—including GPT-4o (Hurst et al. 2024), Intern3-VL (Zhu et al. 2025), Qwen2-Audio (Chu et al. 2023), and Qwen2.5-Omni (Xu et al. 2025) with human standard.
$\mathrm{Avg}_{L1}$ $\mathrm{Avg}_{L2}$ FR1 FR2 LD RR1 RR2 ES PC
Models L1 L2 L3 L4
Baseline 0.410 0.567 0.770 0.740 0.550 0.713 0.650 0.607 0.399
+ Stg.1 0.555 0.548 0.720 0.557 0.540 0.707 0.620 0.593 0.540
+ Stg.1-2 0.554 0.565 0.775 0.687 0.545 0.693 0.625 0.593 0.625
+ Stg.1-3 0.563 0.603 0.780 0.687 0.555 0.720 0.690 0.620 0.619
+ PE - - 0.790 0.763 0.523 0.720 0.625 0.600 0.436
Table 2: Evaluation of Multi-Stage Omni-Modal Reinforcement Learning and Training-Free Prompt Enhancement. We report the average scores for L1 and L2, as well as detailed scores for each high-level task. PE represents Prompt Enhancement.
Figure 4: Modality Ablation Across Different Tasks.We visualize the contributions of different modalities across tasks ranging from perception to interaction. The left 4 tasks are video-based questions, and the right 2 tasks are audio-based questions. Note that ASR-transcribed text (T) was used exclusively for this ablation study, not in Table 1.
Results
Human Level Performance.
Human evaluators achieve an average accuracy of 87.5% on our benchmark, outperforming the best-performing model by a margin of 29.7%. As shown in Figure 3, a substantial performance gap still exists between all models and human-level performance, especially in complex L3-L4 tasks, highlighting a significant need for improvement in the capabilities of current MLLMs on human-centered tasks.
Figure 5: Examples of Successful Reasoning. These cases cover four high-level perception and interaction tasks, including both video-based and audio-based questions. The reasoning processes all demonstrate thinking that integrates characteristics, emotions, and context, and then provides appropriate feedback.
Visual LLMs.
As Table 1, Intern3-VL shows the most advantages in terms of average performance, excelling in both L1-L2 perceptual tasks and L3-L4 high-level tasks. Specifically, LLaVA-OneVision achieves the highest metrics in L1 series tasks, reflecting its exceptional visual perception ability on basic tasks. InternVL3 stands out in L2 and some L3 tasks, demonstrating strong long-video memory and contextual understanding, which is consistent with its outstanding performance in video-related benchmarks (Zhu et al. 2025). Notably, in high-level L3-L4 tasks, all models exhibit metrics ranging from 40 to 60, with little variation, suggesting that visual modality input alone is insufficient to provide adequate information for these tasks.
Omni-models and Audio LLMs.
The inclusion of audio grants omni-models a significant edge over visual-only LLMs in high-level tasks (L3, L4), such as Rapport Recognition and Lie Detection. This cross-modal synergy is further underscored in the Fraud Recognition task, where the Qwen2.5-Omni (0.74) outperforms its specialized audio counterpart. However, this perceptual advantage diminishes in more complex tasks like Psychological Chat. This highlights a limitation for current leading omni-models: the primary bottleneck is not low-level perception but rather a deficiency in high-level, long-context reasoning, which is essential for truly human-centered understanding.
The charts in Figure 3 visually demonstrate the omni-models, particularly Qwen2.5-Omni, show a marked advantage over the visual-only LLM in higher-level tasks, highlighting the critical role of multimodal, including auditory, perception.
Modality Ablation
To conduct a fine-grained analysis of the importance of different modalities, we performed a series of ablation studies using the Qwen2.5-Omni and Qwen2-Audio models. For clarity, we denote video, audio, and the ASR-transcribed text from video as V, A, and T, respectively. Our experiments focused on six challenging tasks from our L3-L4 difficulty tiers, comprising four video-based and two audio-based tasks. We designed two specific experimental settings to probe modal contributions: (1) augmenting both models with ASR-transcribed text (T) as an additional input, and (2) evaluating the Omni-model in a visual-only setting, removing the audio input.
Figure 4 demonstrates the results for our modality ablation study. For video tasks, audio input (A) serves as a powerful supplement to video (V) for the omni-model, significantly enhancing performance on high-level tasks such as rapport and relationship recognition. Incorporating ASR text (T) provides minimal additional benefit, indicating that raw audio already delivers substantial semantic information for this advanced model. In contrast, for audio-based tasks, the audio-only model exhibits a clear dependence on textual input, highlighting its limited speech comprehension and reliance on explicit semantic support. However, the omni-model gains little from text inputs, demonstrating the advantage of comprehensive multi-modal training. Furthermore, all models perform worse on response-related tasks compared to perceptual tasks, highlighting the importance of improving the response capabilities of MLLMs in interactive scenarios.
From Perception to Responses
The above evaluations confirm that visual, auditory, and textual information all play important roles in high-level tasks. Through data observation, we also find that appropriate feedback in communication relies on thorough consideration of omni-modal context and insight into the interlocutor’s needs, emotions, and personal characteristics. Therefore, we believe that reasoning capabilities based on omni-modal inputs are key to enhancing the cognitive and interactive capabilities of MLLMs. In the following sections, we employ a multi-stage, omni-modal reinforcement learning approach to build a reasoning omni model.
Omni-modal Reinforcement Learning
We construct a training set using data independent from the benchmark. We apply Group Relative Policy Optimization (GRPO) (Shao et al. 2024) and design a multi-stage, omni-modal training approach that exposes the Qwen2.5-Omni-7B to all modalities during the Reinforcement Learning with Verifiable Reward (RLVR), enhancing training stability and progressively strengthening perceptual capabilities.
Specifically, in the first stage, we train using pure video frames and textual Question-Answer(QA) pairs to establish reasoning capabilities that integrate visual perception. In the second stage, we train with audio-based QA to develop reasoning that incorporates auditory perception. Finally, we utilize complete video-audio QA to reinforce reasoning that combines both visual and auditory perception. More training details are included in the supplementary materials.
As shown in Table 2, for vision-centric L1 tasks, stage-1 training already yields significant improvements, indicating that reasoning grounded in visual perception can enhance performance on such tasks. For the audio-related tasks PC and FR2, stage-2 training leads to notable gains over stage-1, revealing the success of incorporating auditory reasoning. Most tasks achieve optimal performance after completing all 3 training stages. We sample correctly answered examples and find that the model is indeed capable of deep thinking by integrating characteristics, emotions, and contextual information, as shown in Figure 5.
Training-Free Prompt Enhancement
We observe that the successful reasoning processes elicited by RL training exhibit a consistent thought pattern: perceiving key characteristics, emotion, and context, followed by thinking and then response. Inspired by this, we believe there exists a training-free approach that can improve existing MLLMs’ performance through prompt enhancement. To this end, we design and test the following prompt template, and find that it also boosts performance on high-level tasks, as shown in Table 2.
Conclusion
We introduce the HumanSense benchmark to explore MLLMs’ capabilities in complex human-centered perception and interaction scenarios. We propose that omni-modal reasoning can enhance MLLMs’ performance on such tasks. We aim to inspire the community to recognize the potential of MLLMs in advancing AI interaction experiences.